# A Role of Environmental Complexity on Representation Learning in Deep Reinforcement Learning Agents

### v1

@@ 3.2.2 Clustering spatial heatmaps
Beyond quantifying the consistency of single node spatial representations, we are motivated in understanding what qualitative features nodes contribute to encoding. In order to do so, we demonstrate an effective technique of drawing signal out of the noise by looking at the features that whole populations of nodes encode. First, we take spatial heatmaps from every agent across all $p$ values and 36 checkpoints, then apply k-means clustering with $k=12$. This value of $k$ tends to yield clusters with interpretable and distinct centers. Despite slight variations when running multiple iterations of k-means, the clustering algorithm also consistently settles on the same features. Notably, these clusters partition such that, when multiplying a cluster center by $-1$ and inverting its sign, it approximates another cluster. This phenomenon is demonstrated in Fig 7 which depicts the cluster centers from one run of k-means, each of which can be visualized by a heatmap (the average of all heatmaps belonging to the cluster).
@@
Between different clusters, the ranges of local variances of nodes are mostly comparable, as shown in Fig 8C. We emphasize the relatively large range of local variances shown by the boxplots (See Fig 5 for some pictures of different variance scores). Every cluster contains nodes ranging from locally consistent to noisy. Despite the large variability of nodes inside each cluster, the clusters themselves are consistent when running k-means multiple times. This suggests that k-means effectively captures common features used by agents at different skill levels and in different strategies. Individual nodes may be spatially noisy, but by combining inputs from a layer, these key features may be effectively transmitted through the network.
In Fig 9 we show a few spatial heatmaps alongside the clusters they are classified into (visualized by their centers). These demonstrate how cluster centers can contextualize what heatmaps that would otherwise be as noisy activations. The cluster visually clarifies the most notable features of a heatmap image, suggesting what information a node may be involved in providing to downstream computations in the network. This clustering approach enhances our understanding of spatial encodings by artificial agents, highlighting how their strategies evolve from using place-field-like behavior and tracking goal location, to using the corridor entrance as a key navigation feature. It also provides context to interpret the activations of nodes that might otherwise be dismissed as being noisy and not providing useful information. This kind of unstructured clustering strategy might have broader applicability to interpreting neuron recordings in biological studies, where data is often noisy. Later in Sec 3.4, we give another example of the effectiveness of analyzing populations of nodes in uncovering representations.
@@ 3.3 Landmark Sensitivity Representation
The next representation we consider is landmark sensitivity, specifically measuring how responsive a node is to changes in an available visual landmark. Recall that the shortcut wall is colored purple unlike other white walls in the environment, and it acts as a salient cue when closed off. We hypothesized that having nodes with a robust representation of this landmark should correlate to an agent’s ability to recognize and utilize the open or closed state of the shortcut, and hence correlate with increased performance. This concept is inspired from humans, where the use of distal and proximal cues is important in navigation skills [3, 21].
To quantify a node’s sensitivity to the purple wall landmark, we compare its spatial activity heatmap from 50 closed shortcut episodes with the heatmap generated from 50 open episodes. A large difference between these two indicates high landmark sensitivity. This measure could alternatively be assessed by comparing average node activations between open and closed episodes, but spatial heatmaps offer a clear visual distinction in node responses under open and closed shortcut states, as seen in examples presented in Fig 10. We define a landmark score above $21$ as indicative of clearly distinct heatmaps, such as the two right-most nodes shown in the examples.
@@
Two key observations arise from this representation score. First, as shown in the subplots of Fig 11A and in the combined plot Fig 11B, there is an initial decrease in landmark sensitivity during the early stages in training, particularly for agents trained with $p>0.1$. This period coincides with the same early period of training where spatial representations develop (as seen from the mean local variance in Sec 3.2.1). The dip suggests that early representation formation is dominated by learning basic navigation skills, rather than the development of distinct strategies for open and closed trials. The higher the $p$, the greater the dip in landmark sensitivity, parallel with less exposure to the landmark. On the other hand, the higher the $p$, the more landmark sensitivity is developed by the end of training.
Our second observation focuses on the correlation between shortcut use rates and landmark sensitivity scores, shown in Fig 11C. Towards the end of training, agents tend to converge toward optimal performance (100% shortcut use rate) and consistently develop average landmark sensitivity scores of around 23 (the plotted lines converging at the top right of the plot in Fig 11C) regardless of $p$ value. With even more training, we expect agents from the lower $p$ values of $0.1$ and $0.2$ might also develop towards this optimum. This convergence indicates that the development of landmark sensitivity is tightly coupled with shortcut taking strategies in our environment.
To conclude, we first observed a dip in landmark representation strength early in training, but increased exposure to the landmark (lower $p$) mitigated the severity of this dip. We also noted that the landmark sensitivity is tightly coupled with shortcut use frequency, and higher $p$ agents that learn to use the shortcut more effectively surpass lower $p$ agents in landmark sensitivity scores after training. These insights suggest that while exposure to landmarks aids in developing representations, effectively integrating them into navigation strategies is more influential to building robust representations. These findings will later be contextualized with predictions for human navigators, where we propose that the dynamics of developing landmark knowledge in humans could mirror these dynamics observed in artificial agents.
@@ 3.4 Population level representations and agent “planning”
In this section, we expand our focus from individual node representations to those spanning an entire network layer. We use the same recurrent layer activations as before but analyze the combined activations of all nodes in the layer. Our preliminary exploration involved applying the nonlinear dimensionality reduction technique UMAP [13], reducing the set of T 64-dimensional vectors (where T is the total number of time steps of data collected in all 100 episodes, and $64$ is the number of nodes the layer) to T 2-dimensional vectors that can be visualized. An example is given in Fig 12A.
To explore the structure of thes population-level activations, we initially applied k-means clustering to the complete set of T 2-dimensional activations. We then mapped out the positions of the agent corresponding to each clustered point. Fig 12 gives a visual demonstration of the process. This k-means clustering revealed that in well-trained agents, clusters frequently emerged that corresponded distinctly to different segments of the agent’s trajectory. The clearest separation of clusters in UMAP space distinguished times when the agent was above the corridor from points below, with clusters rarely containing both types of points. Additionally, clusters often formed such that points from trajectories where the agent took the long path and trajectories where the agent took the shortcut were placed in separate clusters (see example Fig 12).
These patterns formed the basis of a population-level representation which we explore further in the following discussion. This approach enables a broader understanding of how collective activations of neural network nodes correlates with an agent’s navigational decisions and strategic “planning” during navigation tasks.
@@ 3.4.1 Quantifying trajectory set separation scores
Figure 12: Example of k-means clustering data points from a $p=0.1$ agent late in training. A. Activations from the recurrent network layer reduced to 2 dimensions via UMAP and clustered according to k-means with $k=4$. B. Each subplot shows where the agent was at time points corresponding to the clustered activation points.
We introduce a population-level representation score, attempting to quantify the idea of network activity separating in activation space based on the parts of a trajectory an agent was on. We first categorize data points from an agent into four sets as shown in Fig 13, noting that these sets are predefined and no longer apply k-means. Positions of the agent for each set are displayed in Fig 13B. From left to right, the first two sets contain points where the agent was below the corridor, and we call them “pre-entrance” (colored orange) if the agent took the longer corridor entrance path on that episode, or “pre-shortcut” (colored green) if the agent took the shortcut. The next two sets contain points where the agent was in the corridor, and similarly are termed “post-entrance” and “post-shortcut”, depending on which path the agent took on that episode. The set definitions are independent of whether the shortcut was actually open on a given episode. These sets aim to capture the “intent” or “plan” of an agent, rather than being based only on its spatial position. Positions in the “pre-entrance” and “pre-shortcut” sets, for example, highly overlap, but their activations in UMAP space can be well separated.
@@
We find that the most interesting set pair is the pre-entrance and pre-shortcut sets (the orange and green points in Fig 13) and is the set separation score that will be the main focus of our analysis. The separation score of this pair increases the most across training and has the best correlation with shortcut usage. In the example shown in Fig 13A, there are some distinct regions that only contain orange points in the top-right and bottom-right of the plot. There is also a region of overlap between the two clusters of interest slightly towards the bottom-right of the image. Notably, these sets separate quite well despite containing many points with similar positions.
@@ 3.4.2 Prescribed trajectories
We interpreted the pre-shortcut and pre-entrance separation score increasing as an indication of a neural network’s growing effectiveness in encoding trajectory or planning information. However, an alternate interpretation might suggest that this increase is not purely due to enhanced encoding abilities, but rather due to changes in policy during training that lead to modified visual inputs. For instance, an agent early in training that only takes shortcuts sometimes will end up with open shortcut episodes in both pre-entrance and pre-shortcut sets. On the other hand, an optimized, well-trained agent might have only open episodes in the pre-shortcut set and closed episodes in the pre-entrance set. The well-trained agent then would have more differing visual stimuli between the two sets, due to the landmark only occurring when the shortcut is closed. One might conclude that an increase in set separation after training is observed due to policy optimization that causes visual stimuli received in the two sets to differ.
@@
To conclusively demonstrate that increased set separation is not only a reflection of changes to behavior, we employ a novel data collection strategy of prescribed trajectories. In this method, we copy the action history of one agent that had approximately $60\%$ shortcut use rate, and then collect activations for all agents under those same actions. This allows us to observe how agents respond to a fixed set of behaviors and visual stimuli. Fig 15B illustrates that, even with controlled actions, agents progressively learn to distinguish between paths leading to the entrance versus the shortcut.
@@
Interestingly, in both alternate measures of set separation shown in Fig 15A and 15B, the difference between different $p$ agents is diminished or eliminated. This suggests that a more optimized policy does contribute to increased set separation score, along with better representation encoding. The key takeaway from this section is that agents learn representations at the population level that indicate an understanding of what trajectory they are headed towards, rather than just their current location. Although closely associated with shortcut usage, these representations also evolve independently of performance, continuing to mature even when policies are nearly optimal. Future research could explore the underlying mechanisms driving this continued representation development, potentially looking into what drives network gradients once earned reward signals have stabilized.
3.4.3 The role of landmark sensitive nodes in trajectory representations
@@
In Sec 3.3, we identified nodes with a landmark representation score greater than $21$ as particularly sensitive to the shortcut wall landmark. Fig 15C demonstrates that these landmark-sensitive nodes contribute more significantly to the separation of pre-shortcut and pre-entrance activations compared to non-sensitive nodes. This analysis only includes open shortcut episodes to minimize the impact of varying visual stimuli on the result.
@@
This finding suggests that even in the absence of the landmark, landmark-sensitive nodes play a crucial role in shaping population-level representations. This could indicate their general importance in advanced navigation or trajectory planning. Alternatively, the landmark sensitivity measure might effectively identifies nodes that consistently provide valuable navigational information to the rest of the network. It is not uncommon to observe some activations in a neural network layer to become less informative over RL training [11, 22], and representation scores may be a way to filter out “useful” nodes. These observations might warrant further study, potentially by examining more quantifiable representations to see if nodes scoring highly on those metrics also play pivotal roles in population codes.
@@
To summarize this whole section, the detailed analysis of trajectory representations implies that agent networks are learning to encode information about “intended” trajectories. Additionally, the correlation between this score with performance is due to more than just differences of visual stimuli between sets. Even though many nodes have a strong spatial component, it is important for us to look at the population activity to derive predictive signals for the agent’s behavior. We have shown examples of population analysis and k-means being useful for discovering ways to interpret noisy activity patterns. We have also shown that population activity as a whole can encode information that is useful in an agent’s policy. This perspective of looking beyond individual nodes to sets could be useful beyond artificial agents, and may have value in understanding collections of spiking neural data, especially where the behavior of individual neurons is noisy and challenging to interpret.
@@ 3.5 Comparisons with human navigators
In this final section, we extrapolate findings from our analysis of artificial agents to human navigators. Our research was inspired by a study examining how the environment individuals live in affect their navigation skills and strategies [1]. Barhorst-Cates et al. assessed individuals from Salt Lake City and Padua on a number of different tasks, including the Dual Solutions Paradigm (DSP), designed to evaluate shortcut usage. Individuals from Padua were found to have a greater ability to use proximal cues during navigation than those from Salt Lake City as well as use shortcuts more often in the DSP.
Drawing parallels, we compare our $p=0.1$ agents to individuals from Salt Lake City, where grid-like street layouts offer few shortcuts, reflected by the sparse shortcut opportunities in the $p=0.1$ environment. On the other hand, Padua’s irregular and organic street layouts are rich in shortcut opportunities and is likened to the $p=0.8$ variation. As shown earlier in Fig 3, the higher the $p$ an agent is trained in, the faster its performance was optimized. Similarly, Barhorst-Cates et al. found individuals from Padua to have generally better navigation skills than those from Salt Lake City. With this comparison, we make two predictions for human navigators and the environments that they live in.

### v2

@@ 3.2.2 Clustering spatial heatmaps
Beyond quantifying the consistency of single node spatial representations, we are motivated in understanding what qualitative features nodes contribute to encoding. In order to do so, we demonstrate an effective technique of drawing signal out of the noise by looking at the features that whole populations of nodes encode. First, we take spatial heatmaps from every agent across all $p$ values and 36 checkpoints, then apply k-means clustering with $k=12$. Despite slight variations when running multiple iterations of k-means, the clustering algorithm consistently settles on similar features. Notably, these clusters partition such that, when multiplying a cluster center by $-1$ and inverting its sign, it approximates another cluster. This phenomenon is demonstrated in Fig 8 which depicts the cluster centers from one run of k-means, each of which can be visualized by a heatmap (the average of all heatmaps belonging to the cluster).
@@
The mean local variance scores of nodes in different clusters are mostly comparable, as shown in Fig 9C. We emphasize that the boxplots show large ranges of mean local variances of constituent nodes (See Fig 6 for some pictures of what heatmaps with different local variance scores look like). Every cluster contains nodes ranging from locally consistent to noisy. Despite this large variability of spatial consistency in each cluster, the features illuminated by the clustering method are consistent across k-means runs. One might imagine that while individual nodes may be spatially noisy, by combining inputs from multiple nodes these features may be effectively transmitted through the network.
To further demonstrate this idea, in Fig 10 we show a few spatial heatmaps alongside the clusters they are classified into (visualized by their centers). By comparing the two, we can qualitatively distinguish the most notable features of a heatmap image, suggesting what information a node may be involved in providing to downstream computations in the network. This clustering approach enhances our understanding of spatial encodings by artificial agents, highlighting how their strategies evolve from using place-field-like behavior and tracking goal location, to using the corridor entrance as a key navigation feature. It also provides context to interpret the activations of nodes that might otherwise be noisy and difficult to interpret. Clustering methods of this sort might have broader applicability to interpreting neuron recordings in biological studies, where data is often noisy, and in general serves as a useful explorative tool to determine what representations are used by agents. Later in Sec 3.4, we give another example of the effectiveness of analyzing populations of nodes in uncovering representations.
@@ 3.3 Landmark Sensitivity Representation
The next representation we consider is landmark sensitivity, specifically measuring how responsive a node is to changes in an available visual landmark. Recall that the shortcut wall is colored pink unlike other white walls in the environment, and it acts as a salient cue when closed off. This leads to an interesting feature of our task environment where cue availability and shortcut opening are coupled together. We hypothesized that having nodes with a robust representation of this landmark should correlate to an agent’s ability to recognize and utilize the open or closed state of the shortcut, and hence correlate with increased performance. This concept is inspired by studies in humans, where the use of distal and proximal cues is important in navigation skills [3, 21]. As we will discuss in this section, landmark encodings are improved by use as part of navigation strategies, and that this frequent exposure alone does lead to development of a robust landmark representation.
To quantify a node’s sensitivity to the pink wall landmark, we compare its spatial activity heatmap from 50 closed shortcut episodes with the heatmap generated from 50 open episodes. A large difference between these two indicates high landmark sensitivity. This measure could alternatively be assessed by comparing average node activations between open and closed episodes, but we use spatial heatmaps which offer a clear visual distinction in node responses under open and closed shortcut states, as seen in examples presented in Fig 11. We note that with a landmark score above around $20$, heatmaps are clearly distinct, such as the two right-most nodes shown in the examples.
@@
Two key observations arise from this landmark sensitivity representation score. First, as shown in the subplots of Fig 12A and in the combined plot Fig 12B, there is an initial decrease in landmark sensitivity during the early stages in training, particularly for agents trained with $p>0.1$. This period coincides with the same early period of training where spatial representations develop (as seen from the mean local variance in Sec 3.2.1). We hypothesize that the natural tendency of neural networks in this task is to first over-emphasize developing location encodings at the cost of other information. Frequent exposure to the landmark however, is preventative for this loss, indicated by a smaller dip in landmark sensitivity the lower $p$ is and the more often and agent sees the landmark. However, we also note that as training progresses, higher $p$ agents that progress in navigation skill more quickly eventually overtake lower $p$ agents in landmark representation scores.
Our second observation focuses on the relationship between shortcut use rates and landmark sensitivity scores, shown in Fig 12C. Towards the end of training, higher p agents tend to converge toward optimal performance (100% shortcut use rate) and consistently develop average landmark sensitivity scores above 20 (the plotted lines converging at the top right of the plot in Fig 12C) regardless of $p$ value. With even more training, we expect agents with lower $p$ values of $0.1$ and $0.2$ might also develop towards this optimum. This convergence indicates that the development of landmark sensitivity is tightly coupled with shortcut taking strategies in our environment.
To conclude, we first observed a dip in landmark representation strength early in training, but increased exposure to the landmark (lower $p$) mitigated the severity of this dip. We also noted that the landmark sensitivity is tightly coupled with shortcut use frequency, and higher $p$ agents that learn to use the shortcut more effectively surpass lower $p$ agents in landmark sensitivity scores after training. These insights suggest that while exposure to landmarks aids in developing representations, effectively integrating them into navigation strategies is more influential to building robust representations than consistent availability of the cue is.
@@ 3.4 Population level representations and agent “planning”
In this section, we expand our focus from individual node representations to those spanning an entire network layer. We use the same recurrent layer activations as before but analyze the combined activations of all nodes in the layer. Our preliminary exploration involved applying the Uniform Manifold Approximation and Projection (UMAP) nonlinear dimensionality reduction technique [13], reducing the set of T 64-dimensional vectors (where T is the total number of time steps of data collected in all 100 evaluation episodes, and $64$ is the number of nodes the layer) to T 2-dimensional vectors that can be visualized. An example is given in Fig 13A. Different points here represent network activities in the recurrent layer at different time steps. UMAP attempts to preserve local distances, so the relative proximity of points indicates similarity of population activities at those time instances.
To explore the structure of the population-level activations, we initially applied k-means clustering to the complete set of T 2-dimensional activations (different clusters represented by different colors in Fig 13A). We then mapped out the positions of the agent corresponding to each clustered point. Fig 13 gives a visual demonstration of the process. This k-means clustering revealed that in well-trained agents, clusters frequently emerged that corresponded distinctly to different segments of the agent’s trajectory. The clearest separation of clusters in UMAP space distinguished times when the agent was above the corridor from points below, with clusters rarely containing both types of points. Additionally, clusters often formed such that points from trajectories where the agent took the long path and trajectories where the agent took the shortcut were placed in separate clusters (see example Fig 13). Note that cluster membership was assigned based only in the similarity of population activity (after UMAP projection). Corresponding similarity of spatial locarions of the agent was found as a consequence.
These patterns formed the basis of a population-level trajectory representation which we explore further in the following discussion. This approach enables a broader understanding of how collective activations of neural network nodes correlates with an agent’s navigational decisions and strategic “planning” during navigation tasks.
@@ 3.4.1 Quantifying trajectory set separation scores
Figure 13: Example of k-means clustering data points from a $p=0.1$ agent late in training. A. Activations from the recurrent network layer reduced to 2 dimensions via UMAP and clustered according to k-means with $k=4$. B. Each subplot shows where the agent was at time points corresponding to the clustered activation points with larger markers demarcating the initial positions of trajectories.
To look further into the relationship between distinct population activation vectors and spatial locations, including those in not-yet-well-trained agents, we introduce a population-level representation score to quantify the idea of network activity separating in activation space based on the parts of a trajectory an agent was on. First, we separate time steps into 4 sets based on the agent’s trajectory Fig 14, following the intuition gained from the previous clustering exploration. Positions of the agent for each set are displayed in Fig 14B. From left to right, the first two sets contain points where the agent was below the corridor, and we call them “pre-entrance” (colored orange) if the agent took the longer corridor entrance path on that episode, or “pre-shortcut” (colored green) if the agent took the shortcut. The next two sets contain points where the agent was in the corridor, and similarly are termed “post-entrance” and “post-shortcut”, depending on which path the agent took on that episode. The set definitions are independent of whether the shortcut was actually open on a given episode. These sets aim to capture the “intent” or “plan” of an agent, rather than being based only on its spatial position. Positions in the “pre-entrance” and “pre-shortcut” sets, for example, have similar spatial points, but their activations in UMAP space can often be separated as seen in the green and orange points in Fig 14A. Note that Fig 14 shows the same set of points as Fig 13 but with grouping based on predetermined trajectory clusters instead of unsupervised clusters.
@@
We find that the most interesting set pair is the pre-entrance and pre-shortcut pair (the orange and green points in Fig 14) and is the set separation score that will be the main focus of our analysis. The separation score of this pair increases the most across training and has the best correlation with shortcut usage. In the example shown in Fig 14A, there are some distinct regions that only contain orange points in the top-right and bottom-right of the plot. There is also a region of overlap between the two clusters of interest slightly towards the bottom-right of the image. In this example, the sets separate quite well despite containing many points with similar positions.
@@ 3.4.2 Prescribed trajectories
We interpreted the pre-shortcut and pre-entrance separation score increasing as an indication of a neural network’s growing effectiveness in encoding trajectory or planning information. However, an alternate interpretation could be that this increase is not purely due to enhanced encoding abilities, but rather due to changes in policy during training that lead to modified visual inputs. For instance, an agent early in training that only takes shortcuts sometimes will end up with open shortcut episodes in both pre-entrance and pre-shortcut sets. On the other hand, an optimized, well-trained agent might have only open episodes in the pre-shortcut set and closed episodes in the pre-entrance set. The well-trained agent then would have more differing visual stimuli between the two sets, due to the landmark only occurring when the shortcut is closed. One might conclude that an increase in set separation after training is observed due to policy optimization that causes visual stimuli received in the two sets to differ.
@@
To conclusively demonstrate that increased set separation is not only a reflection of changes to behavior, we employ a novel network activity collection strategy of prescribed trajectories. In this method, we copy the action history of one agent that had approximately $60\%$ shortcut use rate, and then collect activations for all agents under those same actions. This allows us to observe how agents respond to a fixed set of behaviors and visual stimuli. Fig 16B illustrates that, even with fixed trajectories, agents progressively learn to distinguish between paths leading to the entrance versus the shortcut.
@@
Interestingly, in both alternate measures of set separation shown in Fig 16A and 16B, the difference between different $p$ agents is diminished or eliminated. This suggests that a more optimized policy does contribute to increased set separation score, along with better representation encoding. The key takeaway from this section is that agents learn representations at the population level that are indicative of what trajectory they are headed towards, rather than just their current location. Although closely associated with shortcut usage, these representations also evolve independently of performance, continuing to mature even when policies are nearly optimal. Future research could explore the underlying mechanisms driving this continued representation development, potentially looking into what drives network gradients once earned reward signals have stabilized.
3.4.3 A role of landmark sensitive nodes in trajectory representations
@@
In Sec 3.3, we categorized nodes with a landmark representation score greater than $20$ as being sensitive to the shortcut wall landmark. Fig 16C demonstrates that these landmark-sensitive nodes contribute more significantly to the separation of pre-shortcut and pre-entrance activations compared to non-sensitive nodes. This analysis only includes open shortcut episodes to minimize the impact of varying visual stimuli on the result.
@@
This finding suggests that even in the absence of the landmark, landmark-sensitive nodes play a role in shaping population-level representations. This could indicate their general importance in advanced navigation or trajectory planning. Alternatively, the landmark sensitivity measure might effectively identify nodes that consistently provide valuable navigational information to the rest of the network. It is not uncommon to observe some activations in a neural network layer to become less informative over RL training [11, 22], and representation scores may be a way to filter out “useful” nodes. These observations may warrant further future study, potentially by exploring additional quantifiable representations on individual nodes to see if nodes scoring highly on those metrics also play pivotal roles in population codes.
@@
To summarize Sec 3.4, our detailed analysis of trajectory representations implies that agent networks learn to encode information about intended trajectories. Additionally, the correlation between this score with performance is due to more than just differences of visual stimuli between sets. Even though many individual nodes have a strong spatial component, we can use population activity to derive predictive signals for the agent’s behavior/policy. This perspective of looking beyond individual spatially-tuned nodes to sets could be useful beyond artificial agents, and may have value in understanding collections of spiking neural data, in particular to look for encodings that may be hard to find in individual neurons.
@@ 3.5 Comparisons with human navigators
In this final section, we attempt to derive insights into human navigators based on findings from our analysis of artificial agents. Our research was inspired by a study examining how the environment individuals live in affect their navigation skills and strategies [1]. Barhorst-Cates et al. assessed individuals from Salt Lake City and Padua on a number of different tasks, including the Dual Solutions Paradigm (DSP), designed to evaluate shortcut usage. Individuals from Padua were found to have a greater ability to use proximal cues during navigation than those from Salt Lake City as well as use shortcuts more often in the DSP.
While acknowledging that our toy navigation environment is abstracted from the real world, we compare our $p=0.1$ agents to individuals from Salt Lake City, where grid-like street layouts offer few shortcuts and mountains offer persistent global navigation landmarks. On the other hand, Padua’s irregular and organic street layouts are rich in shortcut opportunities and is likened to the $p=0.8$ variation. This comparison matches learning curves shown earlier in Fig 4 where the higher the $p$ an agent was trained in, the faster its performance developed. Barhorst-Cates et al. similarly found that individuals from Padua to have generally better navigation skills than those from Salt Lake City. With this framework, we make two predictions for human navigators and the environments that they live in.